# MKG-Rank: Enhancing Large Language Models with Knowledge Graph for Multilingual Medical Question Answering

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable progress in medical question answering (QA), yet their effectiveness remains predominantly limited to English due to imbalanced multilingual training data and scarce medical resources for low-resource languages. To address this critical language gap in medical QA, we propose Multilingual Knowledge Graph-based Retrieval Ranking (MKG-Rank), a knowledge graph-enhanced framework that enables English-centric LLMs to perform multilingual medical QA. Through a word-level translation mechanism, our framework efficiently integrates comprehensive English-centric medical knowledge graphs into LLM reasoning at a low cost, mitigating cross-lingual semantic distortion and achieving precise medical QA across language barriers. To enhance efficiency, we introduce caching and multi-angle ranking strategies to optimize the retrieval process, significantly reducing response times and prioritizing relevant medical knowledge. Extensive evaluations on multilingual medical QA benchmarks across Chinese, Japanese, Korean, and Swahili demonstrate that MKG-Rank consistently outperforms zero-shot LLMs, achieving maximum 33.89% increase in accuracy, while maintaining an average retrieval time of only 0.0009 seconds. [^1] Our anonymous code is available at https://anonymous.4open.science/r/MKG-Rank-6B72.
@@ 1 Introduction
Large Language Models (LLMs) Hurst et al. (2024); Anthropic (2024b); Dubey et al. (2024) have achieved remarkable performance in a wide range of Natural Language Processing (NLP) tasks, including question answering and information retrieval Wang et al. (2023); Fan et al. (2024). Beyond general NLP applications, LLMs have also been successfully applied to specific professional domains such as medicine and law, demonstrating promising results Yang et al. (2024c, d); Ke et al. (2024); Zakka et al. (2024); Bernsohn et al. (2024).
While these advances have been achieved predominantly in English-language settings, research on their effectiveness in other languages remains relatively underexplored, especially in medical question-answering Singh et al. (2024a). Specifically, the remaining challenges are: (1) Mainstream LLMs are predominantly trained with English-centric data, resulting in a highly unbalanced distribution between languages Chataigner et al. (2024), which limits their effectiveness in multilingual contexts. (2) High-quality external medical data for low-resource languages are extremely scarce Quercia et al. (2024). As a result, existing LLMs exhibit significant performance gaps in multilingual medical applications, limiting their use in non-English-speaking medical settings.
@@
In summary, our contributions are: (1) We propose MKG-Rank, an efficient framework that enables English-centric LLMs to handle multilingual medical QA by using easily accessible external medical knowledge graphs. (2) We introduce a word-level translation mechanism that ensures precise medical term translation at minimal cost while preventing semantic distortion. (3) We conduct extensive experiments on four multilingual medical QA datasets, showing that MKG-Rank consistently outperforms zero-shot base LLMs, with accuracy improvements of up to 33.89%.
@@ 2 Methodology
In this section, we introduce our proposed MKG-Rank framework, as illustrated in Figure 1. The MKG-Rank framework includes four steps: (1) We first extract medical entities from the question and options and translate them into English. (2) Translated medical entities are used to query an external medical knowledge base (UMLS Bodenreider (2004)) and retrieve relevant knowledge graphs (KGs). (3) We propose a multi-angle ranking strategy to select the most relevant medical triplets. (4) These triplets are first converted into declarative sentences and then, along with the original question and options, fed into the LLM for inference. Additionally, considering the case of invalid external medical knowledge retrieval, we perform self-knowledge mining to acquire relevant knowledge chunks, enhancing LLM’s reasoning in medical QA.
@@ 2.2 External Medical KGs Retrieval
Table 1: Accuracy comparison between our proposed MKG-Rank and the base models on four multilingual datasets: JMMLU (Japanese), CMMLU (Chinese), SW MMLU (Swahili), and KOMMLU (Korean). * marks the base model on which MKG-Rank achieves the greatest performance improvement. Bold indicates the best score for MKG-Rank in each benchmark.
@@ 3.3 Results and Analysis
In Table 1, we compare MKG-Rank (LLM backbone) with the baseline LLMs (zero-shot setting). The results demonstrate that our method consistently outperforms all base LLMs. Specifically, with LLaMA 3.1 70B, we achieved over a 20% improvement across all datasets. For large-scale closed-source LLMs, our method achieved the highest gain on Claude 3.5 Haiku, with an average improvement of 11.1% across the four datasets. Additionally, we achieved an average improvement of 5.53% and 7.87% on GPT-4o-mini and GPT-4o, respectively.
@@ 3.5 Case Study
For scenario 2, we consider cases where the retrieved medical KGs are invalid (e.g., Cylindrical power) due to the question and options containing minimal medical entities. In such cases, we filter out all irrelevant information, aiming to mine the model’s own knowledge to serve as background knowledge for the final reasoning process, thereby helping the LLM make the final decision.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable progress in medical question answering (QA), yet their effectiveness remains predominantly limited to English due to imbalanced multilingual training data and scarce medical resources for low-resource languages. To address this critical language gap in medical QA, we propose Multilingual Knowledge Graph-based Retrieval Ranking (MKG-Rank), a knowledge graph-enhanced framework that enables English-centric LLMs to perform multilingual medical QA. Through a word-level translation mechanism, our framework efficiently integrates comprehensive English-centric medical knowledge graphs into LLM reasoning at a low cost, mitigating cross-lingual semantic distortion and achieving precise medical QA across language barriers. To enhance efficiency, we introduce caching and multi-angle ranking strategies to optimize the retrieval process, significantly reducing response times and prioritizing relevant medical knowledge. Extensive evaluations on multilingual medical QA benchmarks across Chinese, Japanese, Korean, and Swahili demonstrate that MKG-Rank consistently outperforms zero-shot LLMs, achieving maximum 35.03% increase in accuracy, while maintaining an average retrieval time of only 0.0009 seconds. [^1] Our anonymous code is available at https://anonymous.4open.science/r/MKG-Rank-6B72.
@@ 1 Introduction
Large Language Models (LLMs) Hurst et al. (2024); Anthropic (2024b); Dubey et al. (2024) have achieved remarkable performance in a wide range of Natural Language Processing (NLP) tasks, including question answering Jiang et al. (2021); Dong et al. (2022) and information retrieval Wang et al. (2023); Fan et al. (2024). Beyond general NLP applications, LLMs have also been successfully applied to specific professional domains such as medicine and law, demonstrating promising results Yang et al. (2024c, d); Ke et al. (2024); Zakka et al. (2024); Bernsohn et al. (2024).
While these advances have been demonstrated predominantly in English-language settings, research on their effectiveness in other languages remains relatively underexplored, especially in medical question-answering Singh et al. (2024a). Specifically, the remaining challenges are: (1) mainstream LLMs are predominantly trained with English-centric data, resulting in a highly unbalanced distribution between languages Chataigner et al. (2024), which limits their effectiveness in multilingual contexts; and (2) high-quality external medical data for low-resource languages are extremely scarce Quercia et al. (2024). As a result, existing LLMs exhibit significant performance gaps in multilingual medical applications, limiting their use in non-English-speaking medical settings.
@@
In summary, our contributions are: (1) we propose MKG-Rank, an efficient framework that enables English-centric LLMs to handle multilingual medical QA by using easily accessible external medical knowledge graphs; (2) we introduce a word-level translation mechanism that ensures precise medical term translation at minimal cost while preventing semantic distortion; and (3) we conduct extensive experiments on four multilingual medical QA datasets, showing that MKG-Rank consistently outperforms zero-shot base LLMs, with accuracy improvements of up to 35.03%.
@@ 2 Methodology
In this section, we introduce our MKG-Rank, as illustrated in Figure 1. The MKG-Rank framework includes four steps. First, we extract medical entities from the question and options and translate them into English. Then, translated medical entities are used to query an external medical knowledge base (UMLS Bodenreider (2004)) and retrieve relevant knowledge graphs (KGs). Next, we propose a multi-angle ranking strategy to select the most relevant medical triplets. Finally, these triplets are first converted into declarative sentences and then, along with the original question and options, fed into the LLM for inference. Additionally, considering the case of invalid external medical knowledge retrieval, we perform self-knowledge mining to acquire relevant knowledge chunks, enhancing LLM’s reasoning in medical QA.
@@ 2.2 External Medical KGs Retrieval
Table 1: Accuracy comparison between our proposed MKG-Rank and the base models on four multilingual datasets: JMMLU (Japanese), CMMLU (Chinese), SW MMLU (Swahili), and KO MMLU (Korean). * indicates the base model on which MKG-Rank achieves the highest performance gain. The best performance is shown in bold.
@@ 3.3 Results and Analysis
In Table 1, we compare MKG-Rank (LLM backbone) with the baseline LLMs (zero-shot setting). The results demonstrate that our method consistently outperforms all base LLMs. Specifically, with LLaMA 3.1 70B, we achieved over a 20% improvement across all datasets. For large-scale closed-source LLMs, our method achieved the highest gain on Claude 3.5 Haiku, with an average improvement of 11.1% across the four datasets. Additionally, we achieved an average improvement of 5.5% and 7.8% on GPT-4o-mini and GPT-4o, respectively.
@@ 3.5 Case Study
For scenario 2, we consider cases where the retrieved medical KGs are invalid (e.g., Cylindrical power) due to the question and options containing minimal medical entities. In such cases, we filter out all irrelevant information, aiming to mine the model’s knowledge to serve as background knowledge for the final reasoning process, thereby helping the LLM make the final decision.